General Meeting Summary 7/30/20

From Cuawiki
Jump to: navigation, search

PARTICIPANTS: Aaron Brown, Bogdan Wojtsekjhowski, Hamlet Mkrtchyan, David Hamilton, Carlos Munoz-Camacho, Tanja Horn, Steven Lassiter, Rolf Ent, Ho San Ko, Paulo Medeiros, Vardan Tadevosyan, Marie Boer, Vladimir Berdnikov, Brad Sawatzky, Petr Stepanov


FRAME (Carlos)

  • Orsay Lab status back to normal
  • Ongoing work: optical fibers tests
  • defining length, polishing and cleaving procedures - easiest to cleave and no polish (polish scratches, impacts boundaries - see photo)


FAST ELECTRONICS GROUP UPDATE (Chris, Brad S.)

  • Procurements of DAQ/readout related items: all in procurement except for HV cables
  • HV cable discussion with DSG coming up - more detail next time
  • Hall C network infrastructure in good shape
  • planning to install a new 40GB switch during the next down
  • Outstanding procurement items for NPS
  • HV connectors
  • model: go straight to hut, no benefit to go through path
  • Samtec connectors HV rating - check with C. Cuevas
  • Replace cdaql3 as primary file server
  • DAQ fibers to SHMS
  • Firmware and clustering code for crystal readout - people assigned to different projects, should start soon


MECHANICAL/DESIGN (Paulo, Steve L.)

  • Target deck reinforcement out for review
  • Magnet deck reinforcement plates - to be released in the next few months
  • Power supply for magnet - need to further investigate if a suitable one is available
  • have some candidates, e.g. Hall A
  • Notes from Steve: main coil is ~1kA at 115Volts; corrector coils are 520A at 42V


DSG UPDATES (Aaron)

  • Work ongoing
  • Meeting with Brad about HV cables scheduled - will have more detail next time


NPS PROPOSALS

  • Positron DVCS (Carlos)
  • Reviews supportive of case
  • Listing of items required to establish a positron beam
  • Important to to have multi-hall operation for positrons
  • Presentation will be on Wednesday - link to BJ available upon request (contact Carlos)
  • TCS (Marie)
  • Working on preparing a response to the reviewers - should be sent to both theory and experiment reviewers, plus Chair of the PAC and Chair of the User Board
  • Good to include a description of overall trigger design in the slides


NEXT MEETING: THURSDAY 27 AUGUST AT 9:00AM (ET)